OpenStudy (asylum15):

A block of concrete of mass 500kg rests on a horizontal plane. The coefficient of friction between the block and the plane is 0.45. (a) If a horizontal force of 1500N fails to move the block what is the force of friction then acting on the block? (b) Determine the horizontal force necessary to just move the block.

OpenStudy (anonymous):

a) For horizontal plane N=mg - normal force of reaction of the surface on which the block rests. F(friction) = N*mu mu - coefficient of friction b) The horizontal force necessary to just move the block equals in magnitude to the Force of friction you found above but its direction is opposite to that of the force of friction.

OpenStudy (asylum15):

You still around Archie?

OpenStudy (anonymous):

Yes I am

OpenStudy (asylum15):

Can you tell me if I'm wrong? a.) F=Mu*N F=Mu*m*g F=(0.45)(500)(9.81) F=2207.25

OpenStudy (anonymous):

It's correct

OpenStudy (asylum15):

Ok, I'm still confused about b? b = 2207.25 but in a different direction?

OpenStudy (anonymous):

It all depends on your choice of positive direction . You can assign negative direction to the Force of friction and you will get in fact F(friction) = - 2207.25 N And ten F(Just to move) = + 2207.25 N

OpenStudy (anonymous):

then instead of ten sorry

OpenStudy (asylum15):

Its a bit of a trick quesiton no?

OpenStudy (asylum15):

Question*

OpenStudy (anonymous):

Ok let show that on diagram

OpenStudy (anonymous):